# Sequential Randomization Tests Using e-values: Applications for trial monitoring

### v3

@@ 1 Introduction
We discuss potential methods for trial monitoring based on previous efforts. For binary outcomes, our construction implements the i-bet framework with a specific adaptive betting strategy. We present methods with unified notation, validity proofs, and simulation studies demonstrating Type I error control and power characteristics.
@@ 3 Simulation Studies
We evaluated operating characteristics by simulation. For each scenario, we calculated the sample size required for a chi-square test to achieve the target power at $\alpha=0.05$, then ran 5,000 simulated trials at that sample size. We used burn-in = 50 patients and ramp = 100 patients. Control arm event rate was 40% in all scenarios.
@@ 4 Continuous Outcomes
The previous approach treats each patient as a single Bernoulli trial: the outcome (event vs. no event) is observed, and we bet on which arm that patient came from. For continuous endpoints, the logic is the same but the signal is richer. Each patient now contributes a continuous measurement (for example, ventilator-free days, change in biomarker, or a physiologic score), and the betting strategy uses how extreme that value is relative to past data. Therefore, defining wager is slightly more granular.
One may extend this to continuous endpoint. The validity still comes only from randomization: under the null hypothesis, the distribution of the continuous outcome is the same in both arms, so the outcome does not help to predict treatment assignment.
@@ 4.6 Simulation overview
We evaluated this using the same design philosophy as for binary approach. For a given standardized effect size (Cohen’s $d$) and target power, we first computed the fixed-sample size required for a two-sample $t$-test at $\alpha=0.05$. We then simulated trials with that sample size, assigning patients 1:1 to intervention or control, with outcomes drawn from normal distributions of equal variance and means differing by $d$ under the alternative.
@@ 5 Time-to-Event Outcomes
pmlr-v146-grunwald21a developed a safe logrank test using e-values under a proportional hazards model with a prior on the hazard ratio. We attempted to construct a nonparametric approach where validity derives solely from randomization, not from a correctly specified hazard model.
@@ 6 Discussion
The binary implementation is a nonparametric sequential test for randomized trials based on the betting framework for e-values (i-bet pmlr-v177-duan22a). The method requires only that treatment assignment is randomized — no distributional assumptions about outcomes are needed. This makes it a robust complement to model-based analyses. It can be extended to continuous endpoints and time-to-event analyses.
@@ 6.1 Operating characteristics
Simulations demonstrate that the methos properly controlled for Type I error (2–3% vs nominal 5%) while providing approximately 50% power for early stopping in trials designed with 80% power, and 63–66% power in trials designed with 90% power. When early stopping occurs, it happens at roughly half the planned sample size.
Both the continuous and the survival implementations had similar characteristics: Power was traded for continuous Type 1 error control. All simulations showed that Type 1 error was exceptionally well controlled but power was nominally low.
@@ 7.4 Version Control
    Third Version (this version): Added continuous and survival endpoints; text adjustments.

### v4

@@ 1 Introduction
We propose e-RT (e-value Randomized Trial), a method for prospective sequential monitoring of randomized trials. Like i-bet (pmlr-v177-duan22a), e-RT uses betting martingales for inference, but differs in key respects: e-RT monitors sequentially as patients enroll rather than analyzing completed trial data; it requires no covariates or working models; and it uses an adaptive betting fraction derived from running outcome estimates rather than fixed wagers. This yields a method with minimal assumptions suitable for real-time trial monitoring.
@@ 3 Simulation Studies
We evaluated operating characteristics or e-RT by simulation. For each scenario, we calculated the sample size required for a chi-square test to achieve the target power at $\alpha=0.05$, then ran 5,000 simulated trials at that sample size. We used burn-in = 50 patients and ramp = 100 patients. Control arm event rate was 40% in all scenarios.
@@ 4 Continuous Outcomes
The e-RT treats each patient as a single Bernoulli trial: the outcome (event vs. no event) is observed, and we bet on which arm that patient came from. For continuous endpoints, the logic is the same but the signal is richer. Each patient now contributes a continuous measurement (for example, ventilator-free days, change in biomarker, or a physiologic score), and the betting strategy uses how extreme that value is relative to past data. Therefore, defining wager is slightly more granular.
One may extend this to continuous endpoint which we shall call e-RTC. The validity still comes only from randomization: under the null hypothesis, the distribution of the continuous outcome is the same in both arms, so the outcome does not help to predict treatment assignment.
@@ 4.6 Simulation overview
We evaluated e-RTC using the same design philosophy as for binary approach. For a given standardized effect size (Cohen’s $d$) and target power, we first computed the fixed-sample size required for a two-sample $t$-test at $\alpha=0.05$. We then simulated trials with that sample size, assigning patients 1:1 to intervention or control, with outcomes drawn from normal distributions of equal variance and means differing by $d$ under the alternative.
@@ 5 Time-to-Event Outcomes
pmlr-v146-grunwald21a developed a safe logrank test using e-values under a proportional hazards model with a prior on the hazard ratio. We attempted to construct a nonparametric approach where validity derives solely from randomization, not from a correctly specified hazard model. We call this e-survival.
@@ 6 Discussion
The e-Rt is an nonparametric sequential test for randomized trials based on the betting framework for e-values (i-bet pmlr-v177-duan22a). The method requires only that treatment assignment is randomized — no distributional assumptions about outcomes are needed. This makes it a robust complement to model-based analyses. It can be extended to continuous endpoints and time-to-event analyses.
@@ 6.1 Operating characteristics
Simulations demonstrate that the method properly controlled for Type I error (2–3% vs nominal 5%) while providing approximately 50% power for early stopping in trials designed with 80% power, and 63–66% power in trials designed with 90% power. When early stopping occurs, it happens at roughly half the planned sample size.
Both the continuous (e-RTC) and the survival (e-survival) implementations had similar characteristics: Power was traded for continuous Type 1 error control. All simulations showed that Type 1 error was exceptionally well controlled but power was nominally low.
@@ 6.3 Relationship to existing work
The binary e-RT shares the same martingale foundation as i-bet but differs in key respects: it operates prospectively as patients enroll rather than retrospectively on completed data; it requires no covariates or working models; and it uses betting fractions that adapt continuously to running outcome estimates rather than fixed-magnitude wagers guided by covariate-based predictions. This yields a simpler method that may be suitable for real-time trial monitoring.
@@ 7.4 Version Control
    Third Version (Dec 11, 2025): Added continuous and survival endpoints; text adjustments.
  4. 4.
    Forth Version (This version): Correction on signal direction on -RTC. Text adjustments
@@ Appendix A R Code
# Learn direction from past data (NEW)
trt_prev <- treatment[1:(i-1)]
mean_trt <- mean(out_prev[trt_prev == 1])
mean_ctrl <- mean(out_prev[trt_prev == 0])
# Handle edge cases early in trial
if (is.nan(mean_trt) $|$ is.nan(mean_ctrl)) {
direction <- 0
} else {
direction <- sign(mean_trt - mean_ctrl)